If X is Normally distributed with a mean of 3 and a variance of 4, find P(|X−3|>1.6) to 2 decimal places.

The probability is:
Answer #1

P(|X - 3| > 1.6)

= P(|\frac{x-6}{2}|> 1.6/2)

= P(|\frac{x-6}{2}|> 0.8)

= P(|z| > 0.8)

= P(z < -0.8 or z > 0.8)

= P(z < -0.8) + P(z > 0.8)

= 0.2119 + 0.2119

= 0.42